OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (-5, 1). y = -5 x = 1 y = 1 x = -5

OpenStudy (shayaan_mustafa):

Hi ahmadnehela :) How are you? i think you should go with y=1

OpenStudy (anonymous):

y=-5

OpenStudy (shayaan_mustafa):

elly394 Can you elaborate?

OpenStudy (anonymous):

ok now im confused lol

OpenStudy (anonymous):

Never mind it is y=1 :)

OpenStudy (anonymous):

I mixed up x and y

OpenStudy (shayaan_mustafa):

Because in question they are asking for horizontal line passing through the point (-5,1)

OpenStudy (anonymous):

ok can u guys help with one more question

OpenStudy (shayaan_mustafa):

I am 100% sure that y=-5 and x=1 are not the answers.

OpenStudy (shayaan_mustafa):

yes, Sure...

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (shayaan_mustafa):

it is y+1 = 4(x+3)

OpenStudy (anonymous):

I agree

OpenStudy (anonymous):

horizontal lines ALWAYS have the form y=[some number]... it doesn't contain the variable x because the slope is zero.

OpenStudy (anonymous):

thanks but how about this Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (anonymous):

so wait the answer is wrong

OpenStudy (shayaan_mustafa):

which?

OpenStudy (anonymous):

oh... then you should've made a new question post to avoid the confusion... especially on my part..:)

OpenStudy (anonymous):

oh sorry but can anyone help with the new question

OpenStudy (shayaan_mustafa):

OK. It should be y=x-6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Aren't none of them perpendicular? Because the slope should be the negative reciprocal of the original slope, or 1/x, right?

OpenStudy (shayaan_mustafa):

for perpendicular lines, the product of slopes of two lines = -1

OpenStudy (anonymous):

I forgot that the value of x here was 1 so that would work

OpenStudy (anonymous):

I was treating x as an unknown slope basically

OpenStudy (shayaan_mustafa):

let the given slope say m1=-1 (in question's equation) and the slope to find is m2=? (in answer's equation) m1*m2=-1 (-1)*m2=-1 m2=1

OpenStudy (shayaan_mustafa):

So slope of the required line should be +1 Now look at the choices which choices have positive slopes?

OpenStudy (anonymous):

Right but I was being stupid and thinking that -x was the slope, instead of -1. If -x was the slope, then the slope of the perpendicular line would be the negative reciprocal, or 1/x so that the product was equal to -1. It was just a silly mistake :)

OpenStudy (shayaan_mustafa):

yes dear..